The question is:
A right triangle has legs of length 5 inches and 12 inches.
A.) Fine the area of a square drawn on the hypotenuse of the triangle.
B.) What is the length of the hypotenuse?

Answer by ankor@dixie-net.com(15745)

The question is:
A right triangle has legs of length 5 inches and 12 inches.
:
A.) Find the area of a square drawn on the hypotenuse of the triangle.
:
Using pythagorus:
a^2 + b^2 = c^2
:
c^2 would be the area of the square formed by the hypotenuse
:
5^2 + 12^2 = c^2
25 + 144 = 169 sq inches the area of the square formed by the hypotenuse
:
:
B.) What is the length of the hypotenuse?
Hypotenuse = sqrt(169) = 13 inches
: